# A spectral method for multi-view subspace learning using the product of projections

### v1

@@ 1. Introduction
Canonical correlation analysis (Hotelling, 1936) is a time-tested approach for finding associated signals across views, but it cannot distinguish individual signals. To address this limitation, many methods have been developed to explicitly model joint and individual signals (Van Deun et al., 2009; Lock et al., 2013; Zhou et al., 2015; Yang & Michailidis, 2016; Feng et al., 2018; Shu et al., 2019; Gaynanova & Li, 2019; Park & Lock, 2020; Murden et al., 2022; Prothero et al., 2024).
@@ 1.2. Related works
Lock et al. (2013); Zhou et al. (2015); Gaynanova & Li (2019); Park & Lock (2020); Yi et al. (2023); Prothero et al. (2024) consider optimization-based approaches for estimating joint and individual subspaces. While effective, these methods can be computationally slow, particularly with high-dimensional data, and often rely on ad-hoc techniques for rank selection (such as permutation or bi-cross-validation). Feng et al. (2018) and Shu et al. (2019) are spectral methods, like ours, and utilize singular value decomposition to derive joint and individual subspaces. Shu et al. (2019) extends canonical correlation analysis to account for individual components. However, the individual components are restricted to be orthogonal, leading to a different model from ours. Feng et al. (2018) uses the spectrum of the average of projection matrices, which contains information about the principle angles between the two subspace views, to estimate joint and individual components. The method allows for non-orthogonal individual signals and is most similar to ours; however, it has certain limitations in comparison, as we discuss below.
Conceptually, Feng et al. (2018) considers the average of projection matrices, while our method is the first to focus on the product of projection matrices. A key disadvantage of using the average of projection matrices is that its spectrum exhibits a provably smaller gap (by a factor of less than one-half in noiseless settings) between individual and joint components than the product. This results in a substantially worse clustering of singular values in practice, especially in moderate to low signal-to-noise regimes (see e.g, Figure 4 in Appendix A.4). The lack of clustering in the spectrum considered by Feng et al. (2018) makes separating the joint directions from the remaining signals more challenging. See § 4 for more numerical comparison and Appendix A.4 for more technical details.
@@ 2.1. Notation
For $a,b\in\mathbb{R}$, let $a\wedge b=\min(a,b)$, $a\vee b=\max(a,b)$. For a matrix $A\in\mathbb{R}^{n\times p}$, write the compact singular value decomposition (SVD) as $A=U\Sigma V^{T}$, where $\Sigma=\text{diag}(\sigma_{1}(A),\sigma_{2}(A),\dots)$ with the singular values $\sigma_{1}(A)\geq\sigma_{2}(A)\geq\dots>0$ in descending order. We denote $\sigma_{\min}(A)=\sigma_{\operatorname*{rank}(A)}(A)$, $\sigma_{\max}(A)=\sigma_{1}(A)$ as the smallest and largest non-trivial singular values. We denote $\sigma(A)$ to be the collection $\{\sigma_{1}(A),\sigma_{2}(A),\dots,\sigma_{\operatorname*{rank}(A)}(A)\}$. We use $\|A\|_{2}=\sigma_{1}(A)$ to denote the spectral norm, $\|A\|_{F}$ to denote the Frobenius norm and $\operatorname*{col}(A)$ to denote the column space of $A$. We write the subspaces in caligraphic, e.g., we say $\mathcal{A}\subset\mathbb{R}^{n}$ to denote a subspace of $\mathbb{R}^{n}$. For the projection onto a subspace $\mathcal{A}$, we write $P_{\mathcal{A}}\in\mathbb{R}^{n\times n}$. For the projection onto the column space of $A$, we use $P_{\operatorname*{col}(A)}$, which can be defined as $P_{A}=UU^{\top}$. We denote the direct sum of two subspaces $\mathcal{A}$ and $\mathcal{B}$ by $\mathcal{A}\oplus\mathcal{B}$. We adopt the convention to denote the submatrix: $A_{[a:b,c:d]}$ represents the $a$-to-$b$-th row, $c$-to-$d$-th column of matrix $A$; we also use $A_{[a:b,:]}$ and $A_{[:,c:d]}$ to represent $a$-to-$b$-th full rows of $A$ and $c$-to-$d$-th full columns of $A$, respectively.
@@ 2.2. Setup
    $\mathcal{J}\perp\mathcal{I}_{k}$, $\mathcal{N}_{k}\perp\mathcal{O}_{l}$ for $k\neq l$;
@@
    $\bigcap\mathcal{N}_{k}=\{\mathbf{0}\}$ and maximal principle angle between $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ is $<\pi/2$.
@@
3. Our method: Product of Projections Decomposition
Our theoretical insights, which subsequently motivate the proposed estimation approach, rely crucially on the product of projection matrices associated with each view’s subspace. The spectrum of this product corresponds to the cosines of principal angles between the pair of subspaces (Bjoerck & Golub, 1971). In noiseless cases, this spectrum can be clustered to perfectly separate joint and individual subspaces (§ 3.1). In the presence of noise, this spectrum is perturbed, and in § 3.2, we quantify these perturbations with respect to each cluster, leading to explicit conditions for correct estimation of joint rank and quantification of corresponding subspace estimation error. As theoretical perturbation bounds depend on unknown quantities, in § 3.3 we estimate these bounds using rotational bootstrap and random matrix theory, leading to a diagnostic plot that visualizes partitioning of the spectrum into clusters and provides insights into the estimation performance. Throughout, we focus on the case of two views, with Appendix A.3 describing the extension to the larger number of views.
@@ 3.1. Product of projections in the noiseless setting
$P_{\mathcal{I}_{1}}P_{\mathcal{I}_{2}}=(P_{\mathcal{N}_{1}}+P_{\mathcal{O}_{1}})(P_{\mathcal{N}_{2}}+P_{\mathcal{O}_{2}})=P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}=U_{\mathcal{N}_{1}}{U_{\mathcal{N}_{1}}^{\top}U_{\mathcal{N}_{2}}}=\left(U_{\mathcal{N}_{1}}H\right)\Sigma\left(K^{\top}U_{\mathcal{N}_{2}}^{\top}\right).$
@@
where the matrices $\left[U_{\mathcal{J}}\quad U_{\mathcal{N}_{1}}H\right]$ and $\left[U_{\mathcal{J}}\quad U_{\mathcal{N}_{2}}K\right]$ are orthonormal. Thus, the spectrum of $P_{\operatorname*{col}(X_{1})}P_{\operatorname*{col}(X_{2})}$ clusters into two components: a cluster with singular values equal to one and a cluster with singular values in the range $[\sigma_{\min}(\Sigma),\sigma_{\max}(\Sigma)]$. Let $r_{\text{clust1}}$ be the number of singular values in the first cluster (corresponding to singular values one). Then, the joint and individual subspaces can be exactly identified via:
@@
Thus, in the noiseless setting, the spectrum of the product of projections $P_{\operatorname*{col}(X_{1})}P_{\operatorname*{col}(X_{2})}$ (equivalently, the set of cosines of principle angles between $\operatorname*{col}(X_{1})$ and $\operatorname*{col}(X_{2})$) can be divided into two groups. The first group contains values equal to one (equivalently, principle angles 00), whose corresponding singular vectors form a basis for the joint subspace. The second group contains values strictly less than one (equivalently, principle angles strictly greater than 00), reflecting the degree of alignment between the individual subspaces $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$. The gap between these two groups, formally defined as one minus the cosine of the smallest principle angle between $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$, narrows as the alignment between individual subspaces increases.
3.2. Product of projection matrices in a noisy setting
@@
Our main theoretical contribution lies in the precise characterization of perturbations in the spectrum of $\widehat{M}:=P_{\operatorname*{col}(\widehat{X}_{1})}P_{\operatorname*{col}(\widehat{X}_{2})}$. Our result provides general guarantees that hold regardless of the estimation procedure used to approximate the signals $(\widehat{X}_{1},\widehat{X}_{2})$. The quality of the chosen estimation procedure is reflected in our analysis via deterministic quantities $\varepsilon_{1}:=\left\|P_{{\operatorname*{col}(X_{1})}}\left(\Delta_{1}+\Delta_{2}+\Delta_{1}\Delta_{2}\right)P_{{\operatorname*{col}(X_{2})}}\right\|_{2}$ and $\varepsilon_{2}:=\left\|P_{{\operatorname*{col}(X_{1})}}\Delta_{2}+\Delta_{1}P_{{\operatorname*{col}(X_{2})}}+\Delta_{1}\Delta_{2}\right\|_{2}$ with $\Delta_{k}:=P_{{\operatorname*{col}(X_{k})}}-P_{\operatorname*{col}(\widehat{X}_{k})}$. The proof is in the Appendix A.1.
@@
Let $\tau_{\min}:=\sigma_{\min}\left(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right)$ and $\tau_{\max}:=\sigma_{\max}\left(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right)$. If $\operatorname*{rank}(X_{1}),\operatorname*{rank}(X_{2})>0$, then the spectrum of $\widehat{M}$ consists of:
@@
    a remaining group of singular values in $[0,\varepsilon_{2}\wedge 1]$.
@@
This theorem states that in the presence of a signal, the spectrum of the product of projection matrices $\widehat{M}$ organizes into three groups: joint, non-orthogonal individual, and noise directions. The constraint $\varepsilon_{1}<\max\{1-\varepsilon_{2},1-\sigma_{\max}(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}})-\varepsilon_{2}\}$ ensures that the singular values corresponding to the joint structure are distinguishable from other clusters, that is the number of singular values above $1-\varepsilon_{1}$ correctly identifies the joint rank. Similarly, the constraint $\varepsilon_{1}+\varepsilon_{2}<\sigma_{\min}(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}})$ ensures that the group of singular values corresponding to the non-orthogonal individual structure is distinguishable from the noise group. When both constraints hold, the spectrum of $\widehat{M}$ has three non-overlapping clusters corresponding to joint, non-orthogonal individual, and noise directions. The gap between the clusters decreases with larger values of $\varepsilon_{1},\varepsilon_{2}$ and larger alignment of the non-orthogonal individual subspaces ${\mathcal{N}_{1}}$ and ${\mathcal{N}_{2}}$. Figure 1 illustrates these theoretical intervals alongside observed singular values of $\widehat{M}$ and true singular values of $P_{\operatorname*{col}(X_{1})}P_{\operatorname*{col}(X_{2})}$.
@@
The error terms $\varepsilon_{1}$, $\varepsilon_{2}$ depend on the noise level: higher noise leads to larger perturbations $\Delta_{k}$ in estimating each view, and thus larger values of $\varepsilon_{1},\varepsilon_{2}$. Figure 1 shows how a smaller signal-to-noise ratio results in smaller cluster gaps. The estimated subspace dimension $\operatorname*{col}(\widehat{X}_{k})$ also affects the size of $\varepsilon_{1}$, $\varepsilon_{2}$. Underestimating the rank (e.g., $\operatorname*{rank}(\widehat{X}_{1})<\operatorname*{rank}(X_{1})$) can lead to large $\varepsilon_{1}$ and $\varepsilon_{2}$, especially if missing directions are aligned with both $\operatorname*{col}(X_{1})$ and $\operatorname*{col}(X_{2})$. In fact, $\varepsilon_{1}=1$ when a missing direction is part of the joint structure, in which cases the effective number of large singular values corresponding to the joint structure gets smaller than the true joint rank. Overestimating the rank (e.g., $\operatorname*{rank}(\widehat{X}_{1})>\operatorname*{rank}(X_{1})$) can also lead to large errors due to increased dimensionality of $\Delta_{k}$, making the third component of $\varepsilon_{2}$, $\|\Delta_{1}\Delta_{2}\|_{2}$, very large, albeit slight overestimation would still keep those terms small. Thus, overestimation is generally preferred to underestimation. We use these insights in § 3.3, providing practical guidelines on marginal rank estimation.
@@
The clustering of the singular values of $\widehat{M}$ enables a natural estimator for the signal subspaces. The joint subspace dimension is estimated as the number of singular values of $\widehat{M}$ above $1-\varepsilon_{1}$, denoted by $\widehat{r}_{\mathcal{J}}$. According to Theorem 3.1, this estimate is exact if $\varepsilon_{1}<\max\{1-\varepsilon_{2},1-\sigma_{\max}(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}})-\varepsilon_{2}\}$. To estimate the joint subspace itself, a candidate approach is to use the span of the first $\widehat{r}_{\mathcal{J}}$ left (or right) singular vectors of $\widehat{M}$. However, this approach arbitrarily favors one view over the other, insisting that the joint subspace is strictly contained in $\operatorname*{col}(\widehat{X}_{1})$. To avoid this, we symmetrize the product of projection matrices by $\widehat{S}:=\frac{1}{2}(\widehat{M}+\widehat{M}^{\top})$, which effectively averages the estimated joint directions from each view, yielding a more accurate estimate. We estimate the joint and individual subspace structures as
(3) $\widehat{\mathcal{J}}&:=\text{span of the first }\widehat{r}_{\mathcal{J}}\text{ singular vectors of }\widehat{S}=\frac{1}{2}\left(P_{\operatorname*{col}(\widehat{X}_{1})}P_{\operatorname*{col}(\widehat{X}_{2})}+P_{\operatorname*{col}(\widehat{X}_{2})}P_{\operatorname*{col}(\widehat{X}_{1})}\right),\\ \widehat{\mathcal{I}}_{k}&:=\text{span of the first }\operatorname*{rank}(\widehat{X}_{k})-\widehat{r}_{\mathcal{J}}\text{ left singular vectors of }P_{\operatorname*{col}(\widehat{X}_{k})}(I-P_{\widehat{\mathcal{J}}}),\end{aligned}$
and the following theorem characterizes their quality with the proof in Appendix A.1.
@@
Define ${R}_{\mathcal{J}}:=P_{\operatorname*{col}(X_{1})}\Delta_{2}+\Delta_{1}P_{\operatorname*{col}(X_{2})}+\Delta_{1}\Delta_{2}$ and $R_{\mathcal{I}_{k}}:=P_{\mathcal{J}}\Delta_{k}+\Delta_{\mathcal{J}}P_{\operatorname*{col}(X_{k})}+\Delta_{\mathcal{J}}\Delta_{k}$, where $\Delta_{\mathcal{J}}:=P_{\widehat{\mathcal{J}}}-P_{\mathcal{J}}$ and $\Delta_{k}:=P_{{\operatorname*{col}(X_{k})}}-P_{\operatorname*{col}(\widehat{X}_{k})}$. Suppose $\varepsilon_{1}<\max\{1-\varepsilon_{2},1-\sigma_{\max}(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}})-\varepsilon_{2}\}$. Then,
$\dim(\widehat{\mathcal{J}})=\dim(\mathcal{J})\quad\text{and}\quad\|P_{\widehat{\mathcal{J}}}-P_{\mathcal{J}}\|_{2}\leq\frac{\dim(\mathcal{J})^{1/2}\|R_{\mathcal{J}}+R^{\top}_{\mathcal{J}}\|_{2}}{1-\sigma_{\max}(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}})}.$
@@
$\|P_{\widehat{\mathcal{I}}_{k}}-P_{\mathcal{I}_{k}}\|_{2}\leq 2\dim(\mathcal{I}_{k})^{1/2}\|R_{\mathcal{I}_{k}}\|_{2}.$

### v2

@@ 1. Introduction
Canonical correlation analysis (Hotelling, 1936) is a time-tested approach for finding associated signals across views, but it cannot distinguish individual signals. To address this limitation, many methods have been developed to explicitly model joint and individual signals (Van Deun et al., 2009; Lock et al., 2013; Zhou et al., 2015; Yang & Michailidis, 2016; Feng et al., 2018; Shu et al., 2019; Gaynanova & Li, 2019; Park & Lock, 2020; Murden et al., 2022; Prothero et al., 2024; Xiao & Xiao, 2024).
@@ 1.2. Related works
Lock et al. (2013); Zhou et al. (2015); Gaynanova & Li (2019); Park & Lock (2020); Yi et al. (2023); Prothero et al. (2024); Xiao & Xiao (2024) consider optimization-based approaches for estimating joint and individual subspaces. While effective, these methods can be computationally slow, particularly with high-dimensional data, and often rely on ad-hoc techniques for rank selection (such as permutation or bi-cross-validation). Xiao & Xiao (2024) is an exception in that it provides a rigorous theoretical analysis of estimation consistency for its optimization-based approach. However, its methodological framework is based on sparse principal component analysis and on estimating eigenvectors corresponding to features. In contrast, our work focuses on joint and individual subspaces, corresponding to samples, based on spectral methods. Feng et al. (2018), Shu et al. (2019) and Chen et al. (2022) are spectral methods, like ours, and utilize singular value decomposition to derive joint and individual subspaces. Shu et al. (2019) extends canonical correlation analysis to account for individual components. However, the individual components are restricted to be orthogonal, leading to a different model from ours. Feng et al. (2018) uses the spectrum of the average of projection matrices, which contains information about the principle angles between the two subspace views, to estimate joint and individual components. The method allows for non-orthogonal individual signals and is most similar to ours; however, it has certain limitations in comparison, as we discuss below. Chen et al. (2022) consider a different data integration setting where data sets are linked by features rather than matched by samples, using a 2 $\times$ 2 factorial design. Similar to Feng et al. (2018), their method relies on the spectrum of the average of projection matrices for estimation.
Conceptually, Feng et al. (2018) and Chen et al. (2022) consider the average of projection matrices, while our method is the first to focus on the product of projection matrices. A key disadvantage of using the average of projection matrices is that its spectrum exhibits a provably smaller gap (by a factor of less than one-half in noiseless settings) between individual and joint components than the product. This results in a substantially worse clustering of singular values in practice, especially in moderate to low signal-to-noise regimes (see e.g, Figure 4 in Appendix A.6). The lack of clustering in the spectrum considered by Feng et al. (2018) makes separating the joint directions from the remaining signals more challenging. See § 5 for more numerical comparison and Appendix A.6 for more technical details.
@@ 2.1. Notation
For $a,b\in\mathbb{R}$, let $a\wedge b=\min(a,b)$, $a\vee b=\max(a,b)$. For a matrix $A\in\mathbb{R}^{n\times p}$, write the compact singular value decomposition (SVD) as $A=U\Sigma V^{T}$, where $\Sigma=\text{diag}(\sigma_{1}(A),\sigma_{2}(A),\dots)$ with the singular values $\sigma_{1}(A)\geq\sigma_{2}(A)\geq\dots>0$ in descending order. We denote $\sigma_{\min}(A)=\sigma_{\operatorname*{rank}(A)}(A)$, $\sigma_{\max}(A)=\sigma_{1}(A)$ as the smallest and largest non-trivial singular values. We denote $\sigma(A)$ to be the collection $\{\sigma_{1}(A),\sigma_{2}(A),\dots,\sigma_{\operatorname*{rank}(A)}(A)\}$. We use $\|A\|_{2}=\sigma_{1}(A)$ to denote the spectral norm, $\|A\|_{F}$ to denote the Frobenius norm and $\operatorname*{col}(A)$ to denote the column space of $A$. We write the subspaces in calligraphic, e.g., we say $\mathcal{A}\subset\mathbb{R}^{n}$ to denote a subspace of $\mathbb{R}^{n}$. For the projection onto a subspace $\mathcal{A}$, we write $P_{\mathcal{A}}\in\mathbb{R}^{n\times n}$. For the projection onto the column space of $A$, we use $P_{\operatorname*{col}(A)}$, which can be defined as $P_{A}=UU^{\top}$. We denote the direct sum of two subspaces $\mathcal{A}$ and $\mathcal{B}$ by $\mathcal{A}\oplus\mathcal{B}$. We adopt the convention to denote the submatrix: $A_{[a:b,c:d]}$ represents the $a$-to-$b$-th row, $c$-to-$d$-th column of matrix $A$; we also use $A_{[a:b,:]}$ and $A_{[:,c:d]}$ to represent $a$-to-$b$-th full rows of $A$ and $c$-to-$d$-th full columns of $A$, respectively. We use the Bachman-Landau symbols $\mathcal{O}$ to describe the limiting behavior of a function (in terms of sample size and number of features). Further, we use $z\asymp 1$ to express $z=\mathcal{O}(1)$ and $1/z=\mathcal{O}(1)$. Finally, we use $z_{1}\gtrsim z_{2}$ to mean that $z_{1}\geq{c}z_{2}$ for some constant $c>0$.
@@ 2.2. Setup
    $\mathcal{J}\perp\mathcal{I}_{k}$ for all $k\in\{1,2\}$, $\mathcal{N}_{k}\perp\mathcal{O}_{l}$ for all $k,l\in\{1,2\}$;
@@
    $\bigcap\mathcal{N}_{k}=\{\mathbf{0}\}$ and maximal principle angle between $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ is $\pi/2$.
@@
3. A New Perspective: Product of Projections Decomposition
Our theoretical insights, which subsequently motivate the proposed estimation approach, rely crucially on the product of projection matrices associated with each view’s subspace. The spectrum of this product corresponds to the cosines of principal angles between the pair of subspaces (Bjoerck & Golub, 1971). In noiseless cases, this spectrum can be clustered to perfectly separate joint and individual subspaces (§ 3.1). In the presence of noise, this spectrum is perturbed, and in § 3.2, we quantify these perturbations with respect to each cluster, leading to explicit conditions for correct estimation of joint rank and quantification of corresponding subspace estimation error. Throughout, we focus on the case of two views, with Appendix A.5 describing the extension to the larger number of views.
@@ 3.1. Product of projections in the noiseless setting
$P_{\mathcal{I}_{1}}P_{\mathcal{I}_{2}}=(P_{\mathcal{N}_{1}}+P_{\mathcal{O}_{1}})(P_{\mathcal{N}_{2}}+P_{\mathcal{O}_{2}})=P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}=U_{\mathcal{N}_{1}}{U_{\mathcal{N}_{1}}^{\top}U_{\mathcal{N}_{2}}}{U_{\mathcal{N}_{2}}^{\top}}=\left(U_{\mathcal{N}_{1}}H\right)\Sigma\left(K^{\top}U_{\mathcal{N}_{2}}^{\top}\right).$
@@
where the matrices $\left[U_{\mathcal{J}}\quad U_{\mathcal{N}_{1}}H\right]$ and $\left[U_{\mathcal{J}}\quad U_{\mathcal{N}_{2}}K\right]$ have orthonormal columns. Thus, the spectrum of $P_{\operatorname*{col}(X_{1})}P_{\operatorname*{col}(X_{2})}$ clusters into two components: a cluster with singular values equal to one and a cluster with singular values in the range $\sigma_{\min}(\Sigma),\sigma_{\max}(\Sigma)]$. Let $r_{\text{clust1}}$ be the number of singular values in the first cluster (corresponding to singular values one). Then, the joint and individual subspaces can be exactly identified via:
@@
Thus, in the noiseless setting, the spectrum of the product of projections $P_{\operatorname*{col}(X_{1})}P_{\operatorname*{col}(X_{2})}$ (equivalently, the set of cosines of principle angles between $\operatorname*{col}(X_{1})$ and $\operatorname*{col}(X_{2})$) can be divided into two groups. The first group contains values equal to one (equivalently, principle angles 0), whose corresponding singular vectors form a basis for the joint subspace. The second group contains values strictly less than one (equivalently, principle angles strictly greater than 0), reflecting the degree of alignment between the individual subspaces $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$. The gap between these two groups, formally defined as one minus the cosine of the smallest principle angle between $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$, narrows as the alignment between individual subspaces increases.
3.2. Deterministic analysis of the product of projection matrices in a noisy setting
@@
One of our main theoretical contributions lies in the precise characterization of perturbations in the spectrum of $\widehat{M}:=P_{\operatorname*{col}(\widehat{X}_{1})}P_{\operatorname*{col}(\widehat{X}_{2})}$. Our result provides general guarantees that hold regardless of the estimation procedure used to approximate the signals $(\widehat{X}_{1},\widehat{X}_{2})$. The quality of the chosen estimation procedure is reflected in our analysis via deterministic quantities $\varepsilon_{1}:=\left\|P_{{\operatorname*{col}(X_{1})}}\left(\Delta_{1}+\Delta_{2}+\Delta_{1}\Delta_{2}\right)P_{{\operatorname*{col}(X_{2})}}\right\|_{2}$ and $\varepsilon_{2}:=\left\|P_{{\operatorname*{col}(X_{1})}}\Delta_{2}+\Delta_{1}P_{{\operatorname*{col}(X_{2})}}+\Delta_{1}\Delta_{2}\right\|_{2}$ with $\Delta_{k}:=P_{{\operatorname*{col}(X_{k})}}-P_{\operatorname*{col}(\widehat{X}_{k})}$. The proof is in the Appendix A.1.
@@
Let $\tau_{\min}:=\sigma_{\min}(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}})$ and $\tau_{\max}:=\sigma_{\max}(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}})=\left\|P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right\|_{2}$. If $\operatorname*{rank}(X_{1}),\operatorname*{rank}(X_{2})>0$, then the spectrum of $\widehat{M}$ is:
@@
    a remaining group of singular values in $0,\varepsilon_{2}\wedge 1]$.
@@
This theorem states that in the presence of a signal, the spectrum of the product of projection matrices $\widehat{M}$ organizes into three groups: joint, non-orthogonal individual, and noise directions. Here $\varepsilon_{1}$ quantifies the maximal downward perturbation of the singular values associated with the joint signal—i.e., how much true singular values of one may shrink toward zero due to noise. These perturbed singular values fall within $1-\varepsilon_{1}\vee 0,1]$. In contrast, $\varepsilon_{2}$ captures the maximal upward perturbation of null (noise-only) singular values, which may become spuriously large due to random overlap; these lie in $0,\varepsilon_{2}\vee 1]$. When these intervals are disjoint, joint and noise directions can be perfectly separated. More specifically, the constraint $\varepsilon_{1}<\max\{1-\varepsilon_{2},1-\left\|P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right\|_{2}-\varepsilon_{2}\}$ ensures that the singular values corresponding to the joint structure are distinguishable from other clusters, that is the number of singular values above $1-\varepsilon_{1}$ correctly identifies the joint rank. Similarly, the constraint $\varepsilon_{1}+\varepsilon_{2}<\sigma_{\min}(P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}})$ ensures that the group of singular values corresponding to the non-orthogonal individual structure is distinguishable from the noise group. When both constraints hold, the spectrum of $\widehat{M}$ has three non-overlapping clusters corresponding to joint, non-orthogonal individual, and noise directions. The gap between the clusters decreases with larger values of $\varepsilon_{1},\varepsilon_{2}$ and larger alignment of the non-orthogonal individual subspaces $\mathcal{N}_{1}$} and $\mathcal{N}_{2}$}. Figure 1 illustrates these theoretical intervals alongside observed singular values of $\widehat{M}$ and true singular values of $P_{\operatorname*{col}(X_{1})}P_{\operatorname*{col}(X_{2})}$.
@@
The error terms $\varepsilon_{1}$, $\varepsilon_{2}$ depend on the noise level: higher noise leads to larger perturbations $\Delta_{k}$ in estimating each view, and thus larger values of $\varepsilon_{1},\varepsilon_{2}$. Figure 1 shows how a smaller signal-to-noise ratio results in smaller cluster gaps. The estimated subspace dimension $\operatorname*{col}(\widehat{X}_{k})$ also affects the size of $\varepsilon_{1}$, $\varepsilon_{2}$. Underestimating the rank (e.g., $\operatorname*{rank}(\widehat{X}_{1})<\operatorname*{rank}(X_{1})$) can lead to large $\varepsilon_{1}$ and $\varepsilon_{2}$, especially if missing directions are aligned with both $\operatorname*{col}(X_{1})$ and $\operatorname*{col}(X_{2})$. In fact, $\varepsilon_{1}=1$ when a missing direction is part of the joint structure, in which cases the effective number of large singular values corresponding to the joint structure gets smaller than the true joint rank. Overestimating the rank (e.g., $\operatorname*{rank}(\widehat{X}_{1})>\operatorname*{rank}(X_{1})$) can also lead to large errors due to increased dimensionality of $\Delta_{k}$, making the third component of $\varepsilon_{2}$, $\|\Delta_{1}\Delta_{2}\|_{2}$, very large, albeit slight overestimation would still keep those terms small. Thus, overestimation is generally preferred to underestimation. We use these insights in § 4, providing practical guidelines on marginal rank estimation.
@@
The clustering of the singular values of $\widehat{M}$ enables a natural estimator for the signal subspaces. The joint subspace dimension is estimated as the number of singular values of $\widehat{M}$ above $1-\varepsilon_{1}$, denoted by $\widehat{r}_{\mathcal{J}}$. According to Theorem 3.1, this estimate is exact if $\varepsilon_{1}<1-\left\|P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right\|_{2}-\varepsilon_{2}$. To estimate the joint subspace itself, a candidate approach is to use the span of the first $\widehat{r}_{\mathcal{J}}$ left (or right) singular vectors of $\widehat{M}$. However, this approach arbitrarily favors one view over the other, insisting that the joint subspace is strictly contained in $\operatorname*{col}(\widehat{X}_{1})$. To avoid this, we symmetrize the product of projection matrices by $\widehat{S}:=\frac{1}{2}(\widehat{M}+\widehat{M}^{\top})$, which effectively averages the estimated joint directions from each view, yielding a more accurate estimate. We estimate the joint and individual subspace structures as
[…]
We use the symmetrized $\widehat{S}$ to be invariant to the order of views with coinciding left and right singular vectors. We further characterize the quality of these estimates, the proof is in Appendix A.1.
@@
Define ${R}_{\mathcal{J}}:=P_{\operatorname*{col}(X_{1})}\Delta_{2}+\Delta_{1}P_{\operatorname*{col}(X_{2})}+\Delta_{1}\Delta_{2}$ and $R_{\mathcal{I}_{k}}:=P_{\mathcal{J}^{\perp}}\Delta_{k}-\Delta_{\mathcal{J}}P_{\operatorname*{col}(X_{k})}+\Delta_{\mathcal{J}}\Delta_{k}$, where $\Delta_{\mathcal{J}}:=P_{\mathcal{J}}-P_{\widehat{\mathcal{J}}}$ and $\Delta_{k}:=P_{{\operatorname*{col}(X_{k})}}-P_{\operatorname*{col}(\widehat{X}_{k})}$. Suppose $\varepsilon_{1}<1-\left\|P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right\|_{2}-\varepsilon_{2}$. Then,
$\dim(\widehat{\mathcal{J}})=\dim(\mathcal{J})\quad\text{and}\quad\left\|P_{\mathcal{J}}-P_{\widehat{\mathcal{J}}}\right\|_{2}\leq\frac{{}\left\|R_{\mathcal{J}}+R^{\top}_{\mathcal{J}}\right\|_{2}}{1-\left\|P_{\mathcal{N}_{1}}P_{\mathcal{N}_{2}}\right\|_{2}}.$
@@
$\left\|P_{\mathcal{I}_{k}}-P_{\widehat{\mathcal{I}}_{k}}\right\|_{2}\leq{2}\left\|R_{\mathcal{I}_{k}}\right\|_{2}.$